OpenStudy (anonymous):

i need a problem am 12 and the 6th

OpenStudy (bahrom7893):

what problem?

OpenStudy (anonymous):

a math problem

OpenStudy (bahrom7893):

what topic, what class elaborate..

OpenStudy (anonymous):

i dont do online school

OpenStudy (bahrom7893):

http://www.webmath.com/

OpenStudy (anonymous):

can u just give me a problem child

OpenStudy (bahrom7893):

first of all, I'm not a child, second of all, I can give you any problem in the world... third of all WHAT SUBJECT??? okay solve this diff eqn: y"=y

OpenStudy (anonymous):

ur mean

OpenStudy (bahrom7893):

I'm not mean, i'm impatient..

OpenStudy (anonymous):

solve x+26=10

OpenStudy (anonymous):

ok i do ur problem annand hold on

OpenStudy (bahrom7893):

x+y = 40 x-y = 34

OpenStudy (anonymous):

x+26=10 -26 26=16

OpenStudy (anonymous):

sorry i took so long my dad wanted me

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

yeah that the subtract sign

OpenStudy (anonymous):

change keep oppsite

OpenStudy (anonymous):

if am correct

OpenStudy (anonymous):

x=10-26=-16

OpenStudy (anonymous):

ok thanx for giving me a problem

OpenStudy (gg):

haahhaha :D she don't have problem, she need to have it :DDD

OpenStudy (anonymous):

x + 0 = 1

OpenStudy (anonymous):

\[x^2(x^3+x^10)+x7=? \] \[x=7\]

OpenStudy (anonymous):

\[x2(x3+x +10)+x7=?\]

OpenStudy (anonymous):

am not doing that hard stuff

OpenStudy (anonymous):

49(343+10)+823543

OpenStudy (anonymous):

49(353)+823543

OpenStudy (anonymous):

17297+823543

OpenStudy (anonymous):

840840

OpenStudy (anonymous):

ok huh give ma all these hard problem